OpenStudy (anonymous):

f(x)= 1/Sqrt(x) , Find using definition of Derivative F(x)=?

OpenStudy (anonymous):

not really i dont really understand it...its was a homework problem assigned

OpenStudy (amistre64):

the "definition" is:\[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

\[f(x)=\frac{1}{\sqrt x}\Rightarrow f'(x)=\lim_{h\to0}\frac{\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt x}}{h}\]

OpenStudy (anonymous):

From there it is all just simplification

OpenStudy (anonymous):

is this the answer

OpenStudy (amistre64):

ow my eyes!!!

OpenStudy (anonymous):

LOL! My eyes they burn!!!!!!!

OpenStudy (amistre64):

\[\frac{\frac{1}{.x+h}-\frac{1}{.x}}{h}\] \[\frac{\frac{.x-.x+h}{.x.x+h}}{h}\] \[\frac{.x-.x+h}{h({.x.x+h})}\] \[\frac{.x-.x+h}{h({.x.x+h})}*\frac{.x+.x+h}{.x+.x+h}\] \[\frac{x-(x+h)}{h({.x.x+h}){(.x+.x+h)})}\] \[\frac{h}{h({.x.x+h}){(.x+.x+h)})}\] \[\frac{1}{({.x.x+h}){(.x+.x+h)})}\] and when h=0 \[\frac{1}{({.x.x}){(.x+.x})}\] \[\frac{1}{x{2\sqrt{x}}}\] \[\frac{1}{{2x~\sqrt{x}}}\]

OpenStudy (amistre64):

so yes, that looks fine :)

OpenStudy (amistre64):

or simply:\[\frac12x^{-3}\]